the meter was defined in 1963 as the length equal to 1,650,763.73 wavelengths of the orange red radiation emitted by the krypton-86 atom (the meter has since been redefined). What is the wavelength of this transition? What is the frequency?
1
Expert's answer
2020-11-25T02:20:17-0500

1m = 1,650,763.73 × wavelength(λ)

"\lambda" = 1/1,650,763.73

"\lambda" = 6.058 × "10^{-7}"m


c = "\lambda"f

f = c/"\lambda"

f = 3×108/6.058× 10-7

f = 4.95 × 1014 Hz